xx CottonmouthBSSN

CottonmouthBSSN

Lucas Sanches Max Morris

Steven R. Brandt

Date

Abstract

Provides the BSSN equations as generated by the EinsteinEngine.

1 Introduction

The Einstein Engine is a DSL and toolkit for creating Cactus thorns. It can generate complete CarpetX thorns from simple recipes written in Python.

1.1 How to set up EinsteinEngine Create a venv and activate it.

xxxpythonx-mxvenvxvenvx&&x.x./venv/bin/activate
xxx

Install the dependencies.

xxxpythonx-mxpipxinstallx-rxrequirements.txt
xxx

Install EinsteinEngine.

xxxpythonx-mxpipxinstallx.
xxx

1.2 How to generate a thorn

Be sure the venv is activated.

xxx.x./venv/bin/activate
xxx

Typecheck the recipe with MyPy.

xxxmypyxrecipes/Cottonmouth/bssnok.py
xxx

Run the recipe.

     xxxpythonxrecipes/Cottonmouth/bssnok.pyx[--vacuum]x[--fd-orderxN]
     xxx

2 Parameters

By default, the code is generated with matter terms on. To disable them, run using the –vacuum flag.

By default, the code will be generated with 4th order finite difference stencils and 5th order Kreiss-Oliger dissipation. Use the –fd-order flag to switch this to Nth order finite difference stencils and (N+1)th Kreiss-Oliger dissipation.

3 Parameters

11 applyxnewradx 10 Scope: restricted01 BOOLEAN

11 Description: Apply NewRadX boundary conditions

11                    10                01  Default: false

11

11 conformalxfactorxfloor10 Scope: restricted01  REAL

11 Description: The conformal factor W will never be smaller than this value

11  Range  10                01  Default: 1e-06

11 *  10

11

11 dissipationxepsilon 10 Scope: restricted01  REAL

11 Description: The ammount of dissipation to add. Should be in the [0, 1/3[ range

11  Range  10                01  Default: 0.32

11 *  10

11

11 etaxb  10 Scope: restricted01  REAL

11 Description: Standard Gamma driver eta gauge parameter. Must be of order 2 / MxADM

11  Range  10                01  Default: 2.0

11 *  10

11 evolvedxlapsexfloor 10 Scope: restricted01  REAL

11 Description: The evolved lapse will never be smaller than this value

11  Range  10                01  Default: 1e-08

11 *  10

11

11 radpowerxat  10 Scope: restricted01  REAL

11 Description: NewRadX radpower for Atxij

11  Range  10                01  Default: 1.0

11 *  10

11

11 radpowerxchi  10 Scope: restricted01  REAL

11 Description: NewRadX radpower for w

11  Range  10                01  Default: 1.0

11 *  10

11

11 radpowerxconfconnect10 Scope: restricted01  REAL

11 Description: NewRadX radpower Gammax

11  Range  10                01  Default: 1.0

11 *  10

11

11 radpowerxevoxlapse 10 Scope: restricted01  REAL

11 Description: NewRadX radpower for the lapse, alpha

11  Range  10                01  Default: 1.0

11 *  10

11

11 radpowerxevoxshift 10 Scope: restricted01  REAL

11 Description: NewRadX radpower for the shift, betax

11  Range  10                01  Default: 1.0

11 *  10

11

11 radpowerxgt  10 Scope: restricted01  REAL

11 Description: NewRadX radpower for gtxij

11  Range  10                01  Default: 1.0

11 *  10

11

11 radpowerxshiftb  10 Scope: restricted01  REAL

11 Description: NewRadX radpower for the aux. shift vector Bx

11  Range  10                01  Default: 1.0

11 *  10

11

11 radpowerxtrk  10 Scope: restricted01  REAL

11 Description: NewRadX radpower for the trace of Kxij

11  Range  10                01  Default: 1.0

11 *  10

11

4 Interfaces General Implements: cottonmouthbssnok4m

Inherits: admbasex driver odesolvers tmunubasex

Grid Variables

4.0.1 PUBLIC GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  at  10                    01  centering 11 centering={0 0 0}
11                    10  AtDD00  01  compact 11 0
11                    10  AtDD01  01 dimensions 11 3
11                    10  AtDD02  01 distribution 11 DEFAULT
11                    10  AtDD11  01 group type 11 GF
11                    10  AtDD12  01  tags 11 rhs=”CottonmouthBSSNOK4m::Atxrhs” parities={+1 +1 +1 -1 -1 +1 -1 +1 -1 +1 +1 +1 +1 -1 -1 +1 +1 +1}
11                    10  AtDD22  01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  atxrhs  10                    01  centering 11 centering={0 0 0}
11                    10  AtxrhsDD00  01  compact 11 0
11                    10  AtxrhsDD01  01 dimensions 11 3
11                    10  AtxrhsDD02  01 distribution 11 DEFAULT
11                    10  AtxrhsDD11  01 group type 11 GF
11                    10  AtxrhsDD12  01  tags 11 checkpoint=”no” parities={+1 +1 +1 -1 -1 +1 -1 +1 -1 +1 +1 +1 +1 -1 -1 +1 +1 +1}
11                    10  AtxrhsDD22  01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 confconnect 10 01 centering 11 centering={0 0 0}
11                    10 ConfConnectU0 01  compact 11 0
11                    10 ConfConnectU1 01 dimensions 11 3
11                    10 ConfConnectU2 01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 rhs=”CottonmouthBSSNOK4m::ConfConnectxrhs” parities={-1 +1 +1 +1 -1 +1 +1 +1 -1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 confconnectxrhs 10                    01  centering 11 centering={0 0 0}
11                    10 ConfConnectxrhsU0 01  compact 11 0
11                    10 ConfConnectxrhsU1 01 dimensions 11 3
11                    10 ConfConnectxrhsU2 01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no” parities={-1 +1 +1 +1 -1 +1 +1 +1 -1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  deltacons  10                    01  centering 11 centering={0 0 0}
11                    10  DeltaConsU0 01  compact 11 0
11                    10  DeltaConsU1 01 dimensions 11 3
11                    10  DeltaConsU2 01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no” parities={-1 +1 +1 +1 -1 +1 +1 +1 -1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  hamcons  10                    01  centering 11 centering={0 0 0}
11                    10  HamCons  01  compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no” parities={+1 +1 +1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  momcons  10                    01  centering 11 centering={0 0 0}
11                    10  MomConsU0  01  compact 11 0
11                    10  MomConsU1  01 dimensions 11 3
11                    10  MomConsU2  01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no” parities={-1 +1 +1 +1 -1 +1 +1 +1 -1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  evoxlapse  10                    01  centering 11 centering={0 0 0}
11                    10  evoxlapse  01  compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 rhs=”CottonmouthBSSNOK4m::evoxlapsexrhs” parities={+1 +1 +1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  evoxlapsexrhs  10                    01  centering 11 centering={0 0 0}
11                    10  evoxlapsexrhs 01  compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no” parities={+1 +1 +1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  evoxshift  10                    01  centering 11 centering={0 0 0}
11                    10  evoxshiftU0  01  compact 11 0
11                    10  evoxshiftU1  01 dimensions 11 3
11                    10  evoxshiftU2  01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 rhs=”CottonmouthBSSNOK4m::evoxshiftxrhs” parities={-1 +1 +1 +1 -1 +1 +1 +1 -1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 evoxshiftxrhs 10 01 centering 11 centering={0 0 0}
11                    10 evoxshiftxrhsU0 01  compact 11 0
11                    10 evoxshiftxrhsU1 01 dimensions 11 3
11                    10 evoxshiftxrhsU2 01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no” parities={-1 +1 +1 +1 -1 +1 +1 +1 -1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gt  10                    01  centering 11 centering={0 0 0}
11                    10  gtDD00  01  compact 11 0
11                    10  gtDD01  01 dimensions 11 3
11                    10  gtDD02  01 distribution 11 DEFAULT
11                    10  gtDD11  01 group type 11 GF
11                    10  gtDD12  01  tags 11 rhs=”CottonmouthBSSNOK4m::gtxrhs” parities={+1 +1 +1 -1 -1 +1 -1 +1 -1 +1 +1 +1 +1 -1 -1 +1 +1 +1}
11                    10  gtDD22  01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  gtxrhs  10                    01  centering 11 centering={0 0 0}
11                    10  gtxrhsDD00  01  compact 11 0
11                    10  gtxrhsDD01  01 dimensions 11 3
11                    10  gtxrhsDD02  01 distribution 11 DEFAULT
11                    10  gtxrhsDD11  01 group type 11 GF
11                    10  gtxrhsDD12  01  tags 11 checkpoint=”no” parities={+1 +1 +1 -1 -1 +1 -1 +1 -1 +1 +1 +1 +1 -1 -1 +1 +1 +1}
11                    10  gtxrhsDD22  01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  shiftxb  10                    01  centering 11 centering={0 0 0}
11                    10  shiftxBU0  01  compact 11 0
11                    10  shiftxBU1  01 dimensions 11 3
11                    10  shiftxBU2  01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 rhs=”CottonmouthBSSNOK4m::shiftxBxrhs” parities={-1 +1 +1 +1 -1 +1 +1 +1 -1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  shiftxbxrhs  10                    01  centering 11 centering={0 0 0}
11                    10  shiftxBxrhsU0 01  compact 11 0
11                    10  shiftxBxrhsU1 01 dimensions 11 3
11                    10  shiftxBxrhsU2 01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no” parities={-1 +1 +1 +1 -1 +1 +1 +1 -1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  trk  10                    01  centering 11 centering={0 0 0}
11                    10  trK  01  compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 rhs=”CottonmouthBSSNOK4m::trKxrhs” parities={+1 +1 +1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  trkxrhs  10                    01  centering 11 centering={0 0 0}
11                    10  trKxrhs  01  compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no” parities={+1 +1 +1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  w  10                    01  centering 11 centering={0 0 0}
11                    10  w  01  compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 rhs=”CottonmouthBSSNOK4m::wxrhs” parities={+1 +1 +1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  wxrhs  10                    01  centering 11 centering={0 0 0}
11                    10  wxrhs  01  compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no” parities={+1 +1 +1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11

Uses header: newradx.hxx 5 Schedule This section lists all the variables which are assigned storage by thorn Cottonmouth/CottonmouthBSSNOK4m. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11 Always:  11 x
11 At[1]  11 x
11 Atxrhs[1]  11 x
11 evoxshiftxrhs[1]  11 x
11 gt[1]  11 x
11 gtxrhs[1]  11 x
11 shiftxB[1]  11 x
11 shiftxBxrhs[1]  11 x
11 trK[1]  11 x
11 trKxrhs[1]  11 x
11 w[1]  11 x
11 wxrhs[1]  11 x
11 ConfConnect[1]  11 x
11 ConfConnectxrhs[1] 11 x
11 DeltaCons[1]  11 x
11 HamCons[1]  11 x
11 MomCons[1]  11 x
11 evoxlapse[1]  11 x
11 evoxlapsexrhs[1]  11 x
11 evoxshift[1]  11 x

11 x  11 x

11

Scheduled Functions

CottonmouthBSSNOKxInitialGroup x adm2bssnxpt1xgroup x group for function ‘adm2bssnxpt1‘. generated by einsteinengine. x

11 x 11 Language: 11 c
11 x 11 Type:  11 group

11

adm2bssnxpt1xgroup x adm2bssnxpt1 x function ‘adm2bssnxpt1‘ generated by einsteinengine.

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbasex::alp(interior)
11 x 11 x  11 admbasex::betax(everywhere)
11 x 11 x  11 admbasex::betay(everywhere)
11 x 11 x  11 admbasex::betaz(everywhere)
11 x 11 x  11 admbasex::dtbetax(interior)
11 x 11 x  11 admbasex::dtbetay(interior)
11 x 11 x  11 admbasex::dtbetaz(interior)
11 x 11 x  11 admbasex::gxx(interior)
11 x 11 x  11 admbasex::gxy(interior)
11 x 11 x  11 admbasex::gxz(interior)
11 x 11 x  11 admbasex::gyy(interior)
11 x 11 x  11 admbasex::gyz(interior)
11 x 11 x  11 admbasex::gzz(interior)
11 x 11 x  11 admbasex::kxx(interior)
11 x 11 x  11 admbasex::kxy(interior)
11 x 11 x  11 admbasex::kxz(interior)
11 x 11 x  11 admbasex::kyy(interior)
11 x 11 x  11 admbasex::kyz(interior)
11 x 11 x  11 admbasex::kzz(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 atdd00(interior)
11 x 11 x  11 atdd01(interior)
11 x 11 x  11 atdd02(interior)
11 x 11 x  11 atdd11(interior)
11 x 11 x  11 atdd12(interior)
11 x 11 x  11 atdd22(interior)
11 x 11 x  11 evoxlapse(interior)
11 x 11 x  11 evoxshiftu0(interior)
11 x 11 x  11 evoxshiftu1(interior)
11 x 11 x  11 evoxshiftu2(interior)
11 x 11 x  11 gtdd00(interior)
11 x 11 x  11 gtdd01(interior)
11 x 11 x  11 gtdd02(interior)
11 x 11 x  11 gtdd11(interior)
11 x 11 x  11 gtdd12(interior)
11 x 11 x  11 gtdd22(interior)
11 x 11 x  11 shiftxbu0(interior)
11 x 11 x  11 shiftxbu1(interior)
11 x 11 x  11 shiftxbu2(interior)
11 x 11 x  11 trk(interior)
11 x 11 x  11 w(interior)

11

CottonmouthBSSNOKxPostStepGroup x enforcexpt1xgroup x group for function ‘enforcexpt1‘. generated by einsteinengine. x

11 x 11 Language: 11 c
11 x 11 Type:  11 group

11

enforcexpt1xgroup x enforcexpt1 x function ‘enforcexpt1‘ generated by einsteinengine.

11 x 11 Language: 11 c
11 x 11 Reads:  11 evoxlapse(interior)
11 x 11 x  11 gtdd00(interior)
11 x 11 x  11 gtdd01(interior)
11 x 11 x  11 gtdd02(interior)
11 x 11 x  11 gtdd11(interior)
11 x 11 x  11 gtdd12(interior)
11 x 11 x  11 gtdd22(interior)
11 x 11 x  11 w(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 evoxlapse(interior)
11 x 11 x  11 gtdd00(interior)
11 x 11 x  11 gtdd01(interior)
11 x 11 x  11 gtdd02(interior)
11 x 11 x  11 gtdd11(interior)
11 x 11 x  11 gtdd12(interior)
11 x 11 x  11 gtdd22(interior)
11 x 11 x  11 w(interior)

11

CottonmouthBSSNOKxPostStepGroup x enforcexpt2xgroup x group for function ‘enforcexpt2‘. generated by einsteinengine. x

11 x 11 After:  11 enforcexpt1xgroup
11 x 11 Language: 11 c
11 x 11 Type:  11 group

11

enforcexpt2xgroup x enforcexpt2 x function ‘enforcexpt2‘ generated by einsteinengine. x

11 x 11 Language: 11 c
11 x 11 Reads:  11 atdd00(interior)
11 x 11 x  11 atdd01(interior)
11 x 11 x  11 atdd02(interior)
11 x 11 x  11 atdd11(interior)
11 x 11 x  11 atdd12(interior)
11 x 11 x  11 atdd22(interior)
11 x 11 x  11 gtdd00(interior)
11 x 11 x  11 gtdd01(interior)
11 x 11 x  11 gtdd02(interior)
11 x 11 x  11 gtdd11(interior)
11 x 11 x  11 gtdd12(interior)
11 x 11 x  11 gtdd22(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 atdd00(interior)
11 x 11 x  11 atdd01(interior)
11 x 11 x  11 atdd02(interior)
11 x 11 x  11 atdd11(interior)
11 x 11 x  11 atdd12(interior)
11 x 11 x  11 atdd22(interior)

11

CottonmouthBSSNOKxRHSGroup x rhsxgroup x group for function ‘rhs‘. generated by einsteinengine. x

11 x 11 Language: 11 c
11 x 11 Type:  11 group

rhsxgroup x rhs x function ‘rhs‘ generated by einsteinengine.

11 x 11 Language: 11 c
11 x 11 Reads:  11 atdd00(everywhere)
11 x 11 x  11 atdd01(everywhere)
11 x 11 x  11 atdd02(everywhere)
11 x 11 x  11 atdd11(everywhere)
11 x 11 x  11 atdd12(everywhere)
11 x 11 x  11 atdd22(everywhere)
11 x 11 x  11 confconnectu0(everywhere)
11 x 11 x  11 confconnectu1(everywhere)
11 x 11 x  11 confconnectu2(everywhere)
11 x 11 x  11 tmunubasex::ettt(interior)
11 x 11 x  11 tmunubasex::ettx(interior)
11 x 11 x  11 tmunubasex::etty(interior)
11 x 11 x  11 tmunubasex::ettz(interior)
11 x 11 x  11 tmunubasex::etxx(interior)
11 x 11 x  11 tmunubasex::etxy(interior)
11 x 11 x  11 tmunubasex::etxz(interior)
11 x 11 x  11 tmunubasex::etyy(interior)
11 x 11 x  11 tmunubasex::etyz(interior)
11 x 11 x  11 tmunubasex::etzz(interior)
11 x 11 x  11 evoxlapse(everywhere)
11 x 11 x  11 evoxshiftu0(everywhere)
11 x 11 x  11 evoxshiftu1(everywhere)
11 x 11 x  11 evoxshiftu2(everywhere)
11 x 11 x  11 gtdd00(everywhere)
11 x 11 x  11 gtdd01(everywhere)
11 x 11 x  11 gtdd02(everywhere)
11 x 11 x  11 gtdd11(everywhere)
11 x 11 x  11 gtdd12(everywhere)
11 x 11 x  11 gtdd22(everywhere)
11 x 11 x  11 shiftxbu0(everywhere)
11 x 11 x  11 shiftxbu1(everywhere)
11 x 11 x  11 shiftxbu2(everywhere)
11 x 11 x  11 trk(everywhere)
11 x 11 x  11 w(everywhere)
11 x 11 Type:  11 function
11 x 11 Writes: 11 atxrhsdd00(interior)
11 x 11 x  11 atxrhsdd01(interior)
11 x 11 x  11 atxrhsdd02(interior)
11 x 11 x  11 atxrhsdd11(interior)
11 x 11 x  11 atxrhsdd12(interior)
11 x 11 x  11 atxrhsdd22(interior)
11 x 11 x  11 confconnectxrhsu0(interior)
11 x 11 x  11 confconnectxrhsu1(interior)
11 x 11 x  11 confconnectxrhsu2(interior)
11 x 11 x  11 evoxlapsexrhs(interior)
11 x 11 x  11 evoxshiftxrhsu0(interior)
11 x 11 x  11 evoxshiftxrhsu1(interior)
11 x 11 x  11 evoxshiftxrhsu2(interior)
11 x 11 x  11 gtxrhsdd00(interior)
11 x 11 x  11 gtxrhsdd01(interior)
11 x 11 x  11 gtxrhsdd02(interior)
11 x 11 x  11 gtxrhsdd11(interior)
11 x 11 x  11 gtxrhsdd12(interior)
11 x 11 x  11 gtxrhsdd22(interior)
11 x 11 x  11 shiftxbxrhsu0(interior)
11 x 11 x  11 shiftxbxrhsu1(interior)
11 x 11 x  11 shiftxbxrhsu2(interior)
11 x 11 x  11 trkxrhs(interior)
11 x 11 x  11 wxrhs(interior)

11

ODESolversxInitial x cottonmouthbssnokxinitialgroup x bssnok initialization routines x

11 x 11 After: 11 admbasexxpostinitial
11 x 11 Type: 11 group

11

ODESolversxPostStep x cottonmouthbssnokxpoststepgroup x bssnok post-step routines x

11 x 11 Before: 11 admbasexxsetadmvars
11 x 11 Type: 11 group

11

ODESolversxRHS x cottonmouthbssnokxrhsgroup x bssnok equations rhs computation x

11 x 11 Type: 11 group

11

CCTKxANALYSIS x cottonmouthbssnokxanalysisgroup x bssnok analysis routines x

11 x 11 Type: 11 group

11

CottonmouthBSSNOKxInitialGroup (conditional) x adm2bssnxpt2xgroup x group for function ‘adm2bssnxpt2‘. generated by einsteinengine. x

11 x 11 After:  11 adm2bssnxpt1xgroup
11 x 11 Language: 11 c
11 x 11 Type:  11 group

11

bssn2admxgroup x syncxbssn x empty function for explicitly syncing variables as provided in an explicitsyncbatch. syncs: at, confconnect, evoxlapse, evoxshift, gt, shiftxb, trk, w. generated by einsteinengine. x

11 x 11 Before: 11 bssn2adm
11 x 11 Language: 11 c
11 x 11 Sync:  11 at
11 x 11 x  11 confconnect
11 x 11 x  11 evoxlapse
11 x 11 x  11 evoxshift

11 x 11 x  11 gt

11 x 11 x  11 shiftxb

11 x 11 x  11 trk
11 x 11 x  11 w

11 x 11 Type:  11 function

11

adm2bssnxpt2xgroup x syncxbssnxpt2 x empty function for explicitly syncing variables as provided in an explicitsyncbatch. syncs: gt. generated by einsteinengine.

11 x 11 Before: 11 adm2bssnxpt2
11 x 11 Language: 11 c
11 x 11 Sync:  11 gt
11 x 11 Type:  11 function

11

CottonmouthBSSNOKxRHSGroup (conditional) x bssnokxapplyxnewradxxw x function for applying newradx boundary conditions: w -> wxrhs. generated by einsteinengine. x

11 x 11 After:  11 applyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 w(interior)
11 x 11 x  11 wxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 wxrhs(interior)

11

CottonmouthBSSNOKxRHSGroup (conditional) x bssnokxapplyxnewradxxgtxxx x function for applying newradx boundary conditions: gt[l0, l0] -> gtxrhs[l0, l0]. generated by einsteinengine. x

11 x 11 After:  11 applyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 gt(interior)
11 x 11 x  11 gtxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 gtxrhs(interior)

11

CottonmouthBSSNOKxRHSGroup (conditional) x bssnokxapplyxnewradxxgtxxy x function for applying newradx boundary conditions: gt[l0, l1] -> gtxrhs[l0, l1]. generated by einsteinengine. x

11 x 11 After:  11 applyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 gt(interior)
11 x 11 x  11 gtxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 gtxrhs(interior)

11

CottonmouthBSSNOKxRHSGroup (conditional) x bssnokxapplyxnewradxxgtxxz x function for applying newradx boundary conditions: gt[l0, l2] -> gtxrhs[l0, l2]. generated by einsteinengine. x

11 x 11 After:  11 applyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 gt(interior)
11 x 11 x  11 gtxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 gtxrhs(interior)

11

CottonmouthBSSNOKxRHSGroup (conditional) x bssnokxapplyxnewradxxgtxyy x function for applying newradx boundary conditions: gt[l1, l1] -> gtxrhs[l1, l1]. generated by einsteinengine.

11 x 11 After:  11 applyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 gt(interior)
11 x 11 x  11 gtxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 gtxrhs(interior)

11

CottonmouthBSSNOKxRHSGroup (conditional) x bssnokxapplyxnewradxxgtxyz x function for applying newradx boundary conditions: gt[l1, l2] -> gtxrhs[l1, l2]. generated by einsteinengine. x

11 x 11 After:  11 applyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 gt(interior)
11 x 11 x  11 gtxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 gtxrhs(interior)

11

CottonmouthBSSNOKxRHSGroup (conditional) x bssnokxapplyxnewradxxgtxzz x function for applying newradx boundary conditions: gt[l2, l2] -> gtxrhs[l2, l2]. generated by einsteinengine. x

11 x 11 After:  11 applyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 gt(interior)
11 x 11 x  11 gtxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 gtxrhs(interior)

11

CottonmouthBSSNOKxRHSGroup (conditional) x bssnokxapplyxnewradxxat x function for applying newradx boundary conditions: at -> atxrhs. generated by einsteinengine. x

11 x 11 After:  11 applyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 at(interior)
11 x 11 x  11 atxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 atxrhs(interior)

11

adm2bssnxpt2xgroup (conditional) x adm2bssnxpt2 x function ‘adm2bssnxpt2‘ generated by einsteinengine.

11 x 11 Language: 11 c
11 x 11 Reads:  11 gtdd00(everywhere)
11 x 11 x  11 gtdd01(everywhere)
11 x 11 x  11 gtdd02(everywhere)
11 x 11 x  11 gtdd11(everywhere)
11 x 11 x  11 gtdd12(everywhere)
11 x 11 x  11 gtdd22(everywhere)
11 x 11 Type:  11 function
11 x 11 Writes: 11 confconnectu0(interior)
11 x 11 x  11 confconnectu1(interior)
11 x 11 x  11 confconnectu2(interior)

11

CottonmouthBSSNOKxRHSGroup (conditional) x bssnokxapplyxnewradxxtrk x function for applying newradx boundary conditions: trk -> trkxrhs. generated by einsteinengine. x

11 x 11 After:  11 applyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 trk(interior)
11 x 11 x  11 trkxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 trkxrhs(interior)

11

CottonmouthBSSNOKxRHSGroup (conditional) x bssnokxapplyxnewradxxconfconnect x function for applying newradx boundary conditions: confconnect -> confconnectxrhs. generated by einsteinengine. x

11 x 11 After:  11 applyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 confconnect(interior)
11 x 11 x  11 confconnectxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 confconnectxrhs(interior)

11

CottonmouthBSSNOKxRHSGroup (conditional) x bssnokxapplyxnewradxxevoxlapse x function for applying newradx boundary conditions: evoxlapse -> evoxlapsexrhs. generated by einsteinengine. x

11 x 11 After:  11 applyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 evoxlapse(interior)
11 x 11 x  11 evoxlapsexrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 evoxlapsexrhs(interior)

11

CottonmouthBSSNOKxRHSGroup (conditional) x bssnokxapplyxnewradxxevoxshift x function for applying newradx boundary conditions: evoxshift -> evoxshiftxrhs. generated by einsteinengine.

11 x 11 After:  11 applyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 evoxshift(interior)
11 x 11 x  11 evoxshiftxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 evoxshiftxrhs(interior)

11

CottonmouthBSSNOKxRHSGroup (conditional) x bssnokxapplyxnewradxxshiftxb x function for applying newradx boundary conditions: shiftxb -> shiftxbxrhs. generated by einsteinengine. x

11 x 11 After:  11 applyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 shiftxb(interior)
11 x 11 x  11 shiftxbxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 shiftxbxrhs(interior)

11

CCTKxPOSTPOSTINITIAL x initxevolvexcottonmouthbssnok4m x group containing all functions from evolve that should run before step 0. generated by einsteinengine. x

11 x 11 Language: 11 c
11 x 11 Type:  11 group

11

CCTKxPOSTPOSTINITIAL x initxanalysisxcottonmouthbssnok4m x group containing all functions from analysis that should run before step 0. generated by einsteinengine. x

11 x 11 After:  11 initxevolvexcottonmouthbssnok4m
11 x 11 Language: 11 c
11 x 11 Type:  11 group

11

CottonmouthBSSNOKxRHSGroup x applyxdissipationxgroup x group for function ‘applyxdissipation‘. generated by einsteinengine. x

11 x 11 After:  11 rhsxgroup
11 x 11 Language: 11 c
11 x 11 Type:  11 group

11

applyxdissipationxgroup x applyxdissipation x function ‘applyxdissipation‘ generated by einsteinengine.

11 x 11 Language: 11 c
11 x 11 Reads:  11 atdd00(everywhere)
11 x 11 x  11 atdd01(everywhere)
11 x 11 x  11 atdd02(everywhere)
11 x 11 x  11 atdd11(everywhere)
11 x 11 x  11 atdd12(everywhere)
11 x 11 x  11 atdd22(everywhere)
11 x 11 x  11 atxrhsdd00(interior)
11 x 11 x  11 atxrhsdd01(interior)
11 x 11 x  11 atxrhsdd02(interior)
11 x 11 x  11 atxrhsdd11(interior)
11 x 11 x  11 atxrhsdd12(interior)
11 x 11 x  11 atxrhsdd22(interior)
11 x 11 x  11 confconnectu0(everywhere)
11 x 11 x  11 confconnectu1(everywhere)
11 x 11 x  11 confconnectu2(everywhere)
11 x 11 x  11 confconnectxrhsu0(interior)
11 x 11 x  11 confconnectxrhsu1(interior)
11 x 11 x  11 confconnectxrhsu2(interior)
11 x 11 x  11 evoxlapse(everywhere)
11 x 11 x  11 evoxlapsexrhs(interior)
11 x 11 x  11 evoxshiftu0(everywhere)
11 x 11 x  11 evoxshiftu1(everywhere)
11 x 11 x  11 evoxshiftu2(everywhere)
11 x 11 x  11 evoxshiftxrhsu0(interior)
11 x 11 x  11 evoxshiftxrhsu1(interior)
11 x 11 x  11 evoxshiftxrhsu2(interior)
11 x 11 x  11 gtdd00(everywhere)
11 x 11 x  11 gtdd01(everywhere)
11 x 11 x  11 gtdd02(everywhere)
11 x 11 x  11 gtdd11(everywhere)
11 x 11 x  11 gtdd12(everywhere)
11 x 11 x  11 gtdd22(everywhere)
11 x 11 x  11 gtxrhsdd00(interior)
11 x 11 x  11 gtxrhsdd01(interior)
11 x 11 x  11 gtxrhsdd02(interior)
11 x 11 x  11 gtxrhsdd11(interior)
11 x 11 x  11 gtxrhsdd12(interior)
11 x 11 x  11 gtxrhsdd22(interior)
11 x 11 x  11 shiftxbu0(everywhere)
11 x 11 x  11 shiftxbu1(everywhere)
11 x 11 x  11 shiftxbu2(everywhere)
11 x 11 x  11 shiftxbxrhsu0(interior)
11 x 11 x  11 shiftxbxrhsu1(interior)
11 x 11 x  11 shiftxbxrhsu2(interior)
11 x 11 x  11 trk(everywhere)
11 x 11 x  11 trkxrhs(interior)
11 x 11 x  11 w(everywhere)
11 x 11 x  11 wxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 atxrhsdd00(interior)
11 x 11 x  11 atxrhsdd01(interior)
11 x 11 x  11 atxrhsdd02(interior)
11 x 11 x  11 atxrhsdd11(interior)
11 x 11 x  11 atxrhsdd12(interior)
11 x 11 x  11 atxrhsdd22(interior)
11 x 11 x  11 confconnectxrhsu0(interior)
11 x 11 x  11 confconnectxrhsu1(interior)
11 x 11 x  11 confconnectxrhsu2(interior)
11 x 11 x  11 evoxlapsexrhs(interior)
11 x 11 x  11 evoxshiftxrhsu0(interior)
11 x 11 x  11 evoxshiftxrhsu1(interior)
11 x 11 x  11 evoxshiftxrhsu2(interior)
11 x 11 x  11 gtxrhsdd00(interior)
11 x 11 x  11 gtxrhsdd01(interior)
11 x 11 x  11 gtxrhsdd02(interior)
11 x 11 x  11 gtxrhsdd11(interior)
11 x 11 x  11 gtxrhsdd12(interior)
11 x 11 x  11 gtxrhsdd22(interior)
11 x 11 x  11 shiftxbxrhsu0(interior)
11 x 11 x  11 shiftxbxrhsu1(interior)
11 x 11 x  11 shiftxbxrhsu2(interior)
11 x 11 x  11 trkxrhs(interior)
11 x 11 x  11 wxrhs(interior)

CottonmouthBSSNOKxPostStepGroup x bssn2admxgroup x group for function ‘bssn2adm‘. generated by einsteinengine. x

11 x 11 After:  11 enforcexpt2xgroup
11 x 11 Language: 11 c
11 x 11 Type:  11 group

11

bssn2admxgroup x bssn2adm x function ‘bssn2adm‘ generated by einsteinengine. x

11 x 11 Language: 11 c
11 x 11 Reads:  11 atdd00(everywhere)
11 x 11 x  11 atdd01(everywhere)
11 x 11 x  11 atdd02(everywhere)
11 x 11 x  11 atdd11(everywhere)
11 x 11 x  11 atdd12(everywhere)
11 x 11 x  11 atdd22(everywhere)
11 x 11 x  11 evoxlapse(everywhere)
11 x 11 x  11 evoxshiftu0(everywhere)
11 x 11 x  11 evoxshiftu1(everywhere)
11 x 11 x  11 evoxshiftu2(everywhere)
11 x 11 x  11 gtdd00(everywhere)
11 x 11 x  11 gtdd01(everywhere)
11 x 11 x  11 gtdd02(everywhere)
11 x 11 x  11 gtdd11(everywhere)
11 x 11 x  11 gtdd12(everywhere)
11 x 11 x  11 gtdd22(everywhere)
11 x 11 x  11 trk(everywhere)
11 x 11 x  11 w(everywhere)
11 x 11 Type:  11 function
11 x 11 Writes: 11 admbasex::alp(everywhere)
11 x 11 x  11 admbasex::betax(everywhere)
11 x 11 x  11 admbasex::betay(everywhere)
11 x 11 x  11 admbasex::betaz(everywhere)
11 x 11 x  11 admbasex::gxx(everywhere)
11 x 11 x  11 admbasex::gxy(everywhere)
11 x 11 x  11 admbasex::gxz(everywhere)
11 x 11 x  11 admbasex::gyy(everywhere)
11 x 11 x  11 admbasex::gyz(everywhere)
11 x 11 x  11 admbasex::gzz(everywhere)
11 x 11 x  11 admbasex::kxx(everywhere)
11 x 11 x  11 admbasex::kxy(everywhere)
11 x 11 x  11 admbasex::kxz(everywhere)
11 x 11 x  11 admbasex::kyy(everywhere)
11 x 11 x  11 admbasex::kyz(everywhere)
11 x 11 x  11 admbasex::kzz(everywhere)

11

CottonmouthBSSNOKxAnalysisGroup x constraintsxgroup x group for function ‘constraints‘. generated by einsteinengine. x

11 x 11 Language: 11 c
11 x 11 Type:  11 group

11

constraintsxgroup x constraints x function ‘constraints‘ generated by einsteinengine. x

11 x 11 Language: 11 c
11 x 11 Reads:  11 atdd00(everywhere)
11 x 11 x  11 atdd01(everywhere)
11 x 11 x  11 atdd02(everywhere)
11 x 11 x  11 atdd11(everywhere)
11 x 11 x  11 atdd12(everywhere)
11 x 11 x  11 atdd22(everywhere)
11 x 11 x  11 confconnectu0(everywhere)
11 x 11 x  11 confconnectu1(everywhere)
11 x 11 x  11 confconnectu2(everywhere)
11 x 11 x  11 tmunubasex::ettt(interior)
11 x 11 x  11 tmunubasex::ettx(interior)
11 x 11 x  11 tmunubasex::etty(interior)
11 x 11 x  11 tmunubasex::ettz(interior)
11 x 11 x  11 tmunubasex::etxx(interior)
11 x 11 x  11 tmunubasex::etxy(interior)
11 x 11 x  11 tmunubasex::etxz(interior)
11 x 11 x  11 tmunubasex::etyy(interior)
11 x 11 x  11 tmunubasex::etyz(interior)
11 x 11 x  11 tmunubasex::etzz(interior)
11 x 11 x  11 evoxlapse(interior)
11 x 11 x  11 evoxshiftu0(interior)
11 x 11 x  11 evoxshiftu1(interior)
11 x 11 x  11 evoxshiftu2(interior)
11 x 11 x  11 gtdd00(everywhere)
11 x 11 x  11 gtdd01(everywhere)
11 x 11 x  11 gtdd02(everywhere)
11 x 11 x  11 gtdd11(everywhere)
11 x 11 x  11 gtdd12(everywhere)
11 x 11 x  11 gtdd22(everywhere)
11 x 11 x  11 trk(everywhere)
11 x 11 x  11 w(everywhere)
11 x 11 Type:  11 function
11 x 11 Writes: 11 deltaconsu0(interior)
11 x 11 x  11 deltaconsu1(interior)
11 x 11 x  11 deltaconsu2(interior)
11 x 11 x  11 hamcons(interior)
11 x 11 x  11 momconsu0(interior)
11 x 11 x  11 momconsu1(interior)
11 x 11 x  11 momconsu2(interior)